OpenStudy (anonymous):

PLEASE HELP AND I WILL GIVE MEDAL! :P For questions 1-5, choose the counterexample that disproves each conjecture. For all real numbers n, 1 ÷ n > 0. n = 1 n = 3 n = 0.5 n = -0.5 Question 2. 2. If n is a prime number, then n2 has a 1, 5, or 9 in the ones place. n = 3 n = 17 n = 31 n = 2

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

If you could help that would be great :)

OpenStudy (aum):

Question 1: Choose n = -0.5. Then, 1 ÷ n = 1 ÷ (-0.5) = -2 which is < 0 disproving the conjecture.

OpenStudy (aum):

Question 2: Choose n = 2. Then, n^2 = 4 disproving the conjecture that if n is a prime number, then n^2 has a 1, 5, or 9 in the ones place.

OpenStudy (anonymous):

Thanks!

OpenStudy (aum):

yw.

OpenStudy (anonymous):

Do you think you could help me with a few more? PLease

OpenStudy (anonymous):

A student concludes that if x is a real number, then x2 > x. x = -2 x = 3 x = 1 x = -1 Question 4. 4. A student concludes that if x is a real number, then x2 _< x3. Option A: ½ Option B: 0 Option C: 1 Option D:3/2 Question 5. 5. After completing several multiplication problems, a student concludes that the product of two binomials is always a trinomial. (x + 1)(x + 3) (x + 2)(x - 2) (x + 4)(y -3) (x + 5)(x - 6)

OpenStudy (anonymous):

@aum